An arrow is launched at a velocity of 20 m/s in a direction making an angle of 25° upward with the horizontal. (a) How long will it take to hit the ground? (b) How far from the cannon ball will it hit the ground? (c) Compute for maximum height to be reached by the cannon ball. 


1
Expert's answer
2021-10-03T13:23:47-0400

(a) The time is


"t=\frac{2v\sin\theta}{g}=1.72\text s."

(b) The range is


"R=\frac{v^2\sin(2\theta)}{g}=31.3\text m."

(c) The maximum height is


"h=\frac{v^2\sin^2\theta}{2g}=3.65\text{m}."